OpenStudy (anonymous):

2. (1 pt) Solve. 3y + 15 > –6

pooja195 (pooja195):

Start by subtracting 15 from both sides 3y + 15 > –6 -6-15=?

OpenStudy (anonymous):

-21

pooja195 (pooja195):

3y + 15 > –6 3y > –21 divide both sides by 3 y>?

OpenStudy (anonymous):

and then 3y - 15 = -13y?

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

?

pooja195 (pooja195):

We are isolating y correct?

OpenStudy (anonymous):

it shows that we are trying to see what y is greater than

OpenStudy (anonymous):

A. y > 3 B. y > –7 C. y > 13 D. y > –17

pooja195 (pooja195):

Correct so basically we are trying to isolate y 3y + 15 > –6 3y > –21 divide both sides by 3 y>? -21/3=?

OpenStudy (anonymous):

7 or -7

OpenStudy (anonymous):

i say 7

pooja195 (pooja195):

Nope we are dividing a negative by a positive the solution will be negative

OpenStudy (anonymous):

k

OpenStudy (anonymous):

so the answer is b

pooja195 (pooja195):

Yes =) Welcome to openstudy! :)

OpenStudy (anonymous):

thx